OpenStudy (anonymous):

6x=3y= 27 change to slope intercept form

OpenStudy (anonymous):

try to arrange it in the form y =mx +c ( that is y with coefficent as +1 should be on 1 side of the equation and the rest on the other side )

OpenStudy (anonymous):

and by the way check ur question ?

OpenStudy (anonymous):

6x+3y= 27 * lol

OpenStudy (anonymous):

3y=-6x+27 y=(-6/3 ) x +27/3 y=-2x+9 ( which is of the form y=mx+c ) here slope(m)=-2 and y-intercept(c) is 9